OpenStudy (anonymous):

Which of the following expressions are equivalent? Justify your reasoning. A. 4√x3 B. 1/(x^-1) C. 10√x5•x4•x2 D. x^1/3* x^1/3* x^1/3

OpenStudy (danjs):

for D, when you multiply powers of the same bases together, add the powers together x^a * x^b * x ^c = x^(a + b + c)

OpenStudy (danjs):

For B) When you move a base to a power to the other side of the fraction, change the sign on the power

OpenStudy (anonymous):

so x^3/3 ?

OpenStudy (danjs):

1 / x^(-a) = x^(+a)

OpenStudy (danjs):

yes, what is 3/3

OpenStudy (anonymous):

B and D are both X, correct?

OpenStudy (danjs):

yep

OpenStudy (anonymous):

so thats the answer? just B and D?

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

umm

OpenStudy (danjs):

A is X^(3/4) and C has a 10 coefficient, so those are not the same

OpenStudy (danjs):

as any other